# Flight Demonstration and Model Validation of a Prototype Variable-Altitude Venus Aerobot

### v1

@@ Abstract
This paper details a significant milestone towards maturing a buoyant aerial robotic platform, or aerobot, for flight in the Venus clouds. We describe two flights of our subscale altitude-controlled aerobot, fabricated from the materials necessary to survive Venus conditions. During these flights over the Nevada Black Rock desert, the prototype flew at the identical atmospheric densities as 54 to 55 km cloud layer altitudes on Venus. We further describe a first-principle aerobot dynamics model which we validate against the Nevada flight data and subsequently employ to predict the performance of future aerobots on Venus. The aerobot discussed in this paper is under JPL development for an in-situ mission flying multiple circumnavigations of Venus, sampling the chemical and physical properties of the planet’s atmosphere and also remotely sensing surface properties.
©2024 California Institute of Technology. Preprint submitted to AIAA Journal of Aircraft.
@@
$\theta$  = angle tangent to ZP balloon at $s$, radians
@@ 1 Introduction
Figure 1: Subscale prototype of a JPL’s variable-altitude controlled aerobot, made from Venus-compatible materials, in flight over the Black Rock desert in Nevada, USA. Photograph taken July 2022.
@@ 1.1 Venus Aerobot Background
Balloon experiments have operated on Venus before but nothing of the proposed capability. The Soviet Vega lander missions of 1985 included balloon probes with two small 7 kg gondolas that flew for approximately 48 hours in the cloud layer at a nominally constant altitude [1]. The challenge today is expand this early demonstrator to a larger and fundamentally more capable buoyant aerial robotic platform (aerobot), thereby enabling the greater science return promised by a dedicated mission. NASA’s Jet Propulsion Lab (JPL) has been developing large aerobots for specifically this purpose for over twenty years, including multiple Discovery proposals [2], carrying a 40-100 kg gondola below the aerobot with science instrumentation consistent with a much more complete characterization of the Venus cloud layer. These aerobots, like the Vega balloons before, were designed to maintain altitude through the use of "superpressure" - i.e. a high strength envelope inflated to pressure slightly above the ambient atmospheric pressure. Given its constant volume, a superpressure balloon loses buoyancy when displaced upwards and gains buoyancy when displaced downwards, making it altitude stable.
@@
More recently, a NASA study described in Cutts et al. [3] refocused JPL efforts on an aerobot that can instead actively control its altitude over a variable range (Figure 1) instead of staying at a constant altitude. A variable-altitude approach allows the reuse of a single payload over a wide range of conditions in the clouds, thereby more fully characterizing the cloud layer (and providing an even greater science return) for a modest complexity cost. The success of the Google Loon program [4] provided credence to the rapid development of variable altitude aerobots for Earth, and the NASA study prioritized the adaption of this technology for Venus. A number of variable-altitude Venus design points and their development are described in [5], with targeted altitude range on Venus from 52 to 62 km and a flight duration goal of 117 Earth-days (one Venus day). Such platforms are highly scaleable, with designs able to carry from 15 kg to 230 kg depending on mission ambition.
@@
This paper describes a significant milestone in the technology advancement of this Venus aerial plaform: JPL’s first outdoor flights of a prototype aerobot, made from Venus-compatible materials, at atmospheric densities within the proposed flight envelope for Venus. Table 2 describes the demonstrated flight parameters of flight test in comparison to Venus. Balloon flight dynamics are principally driven by atmospheric density, meaning that this flight recreates a relevant environment for flight dynamics in Venus’s cloud layer from 54-55 km above the surface. The targeted Venus temperatures are roughly $15^{\circ}$ greater than those in Nevada however, so models must be employed to take this and other environmental differences into account. Accordingly, a key objective for this flight was to provide validation data for our flight dynamics model FLOATS (FLight Operations and Aerobot Trajectory Simulator), which we subsequently use to predict mission trajectories and operations on Venus.
@@
(incl. $+20^{\circ}$ local offset condition) [7]
@@
Table 2: Atmospheric equivalencies between Venus and test flight in Nevada, USA. Altitudes flown highlighted in blue. Note exact match in atmospheric density, although equivalent densities at Venus are slightly warmer. While flights at higher altitudes are possible with our platform, balloons must trade altitude for payload mass for a given envelope volume, and a complete characterization of the balloon dynamics for FLOATS required more payload mass than the full altitude range could support on a prototype of this scale.
@@ 1.2 Programmatic Context
The variable-altitude aerobot was accordingly adopted into two OWL-support mission concept studies for Venus [10, 11], and Cutts et al. [12] describe a variety of smaller missions that could be performed with a variable-altitude aerobot, including concepts for New Frontiers and Discovery proposals. The prototype flown in the flight tests described in this paper is approximately one-third scale of the favored design point of these previous mission studies.
@@ 1.3 Paper Outline
A video of the test flights in this paper can be found at https://www.jpl.nasa.gov/news/jpls-venus-aerial-robotic-balloon-prototype-aces-test-flights.
@@ 2.1 Aerobot Prototype Specifications
Figure 2: Photo of aerobot at takeoff, with the two balloon chambers labeled. The internal SP balloon acts as a helium reservoir, and the remaining helium rests in the outer ZP balloon. Both balloons are made from the materials needed for a Venus mission, and helium is pumped/vented between balloons for altitude control.
@@
Figure 3 illustrates the design of the two balloon envelopes. The outer ZP is a bilaminate of Teflon® fluorinated ethylene propylene (FEP) and Kapton® polyimide, with a silver metalization layer and bonding adhesive. The Teflon® offers excellent resistance to the clouds of sulfuric acid aerosols during Venus environment evaluations [18], while the Kapton® provides a helium gas barrier and the tensile strength for carrying the payload. The metalization reduces the effect of helium diffusion through the material and further acts as a second surface mirror (with a measured solar absorptivity $\alpha=0.08$ and infrared emissivity $\epsilon=0.52$) to reduce solar heating of the helium. The bilaminate design further adds resistance to pinholing from manufacturing defects, as the chances of two defects aligning between the layers is extremely small. The ZP balloon is constructed as a 60-degree sphere-cone-sphere with upper diameter of 5 meters and lower diameter of 2.5 meters.
Figure 3: Aerobot balloon envelope designs. A: Material stackup of the zero-pressure (ZP) balloon, consisting of a metalized FEP-polyimide bilaminate. B: Two layers of the super-pressure (SP) balloon envelope, consisting of a Vectran fabric with internal urethane gas barrier. C: Photograph of inflated SP balloon. Reproduced with permission from Izraelevitz et al. [13]
The inner SP balloon is fabricated as a 2.5 m diameter sphere, also two-layered, with a Vectran® fabric (200 denier 2x2 basket weave at 76 threads-per-inch) to provide pressurization strength and a heat-sealed thermoplastic urethane bladder for gas retention. The two layers are indexed together with attachment tabs to keep relative alignment. The SP balloon was proof-tested to 14 kPa before the flight, and an identical SP balloon was taken to 61 kPa to provide confidence in the planned pressurization margin utilized in the flight tests.
@@ 2.2 Instrumentation, Gas Transfer, and Communications Payload
For accurate measurement of the flight dynamics of the aerobot, we designed and implemented an instrumentation system (Figure 4) that could not only gather information about the 6-DOF dynamics of the aerobot, but also on its thermal state and driving inputs from the environment such as atmospheric temperature, pressure, wind, and solar radiation. In addition, our instrumentation package coordinated the gas transfer between the two aerobot chambers, executed flight termination processes, and sent back real-time telemetry. This payload module was fabricated using hobby-grade avionics and commercial-of-the-shelf instrumentation for ease of development timeline, and would be replaced by spaceflight-grade systems for a Venus mission. This payload is a second iteration of what was flown in our previous indoor flights in Tillamook Airship Hangar [13] with additions for wind measurements, solar forcing, long distance commanding, and Federal Aviation Administration requirements. Components are summarized in Table 3.
@@
InertialSense $\mu$INS Inertial Navigation System (INS) with dual GPS receiver on 1m baseline BCM, Gondola, Ground 30
Texas Instruments LM35-CA/Z Skin temperature sensor x4 BCM 20
@@
Apogee Instruments SP510/610 Thermopile pyranometers (solar radiation, up to 4 $\mu$m wavelength) x4 Gondola, Ground 30
Apogee Instruments SL510-SS/610-SS Pyrgeometer (IR radiation 4–100 $\mu$m) x2 Gondola, Ground 30
Windmaster Pro Ultrasonic anemometer for relative wind Gondola 20
@@ 2.2.1 Payload Platforms
    Buoyancy Control Module (BCM): The BCM contained the entire gas transfer system, including a 12.7 mm (1/2 inch) inner diameter bypass tube through the entire vertical extent of the SP balloon to transfer gas into the ZP balloon. The BCM system needed to be as close to the balloon chambers as possible to reduce additional flow path length and associated loss of pressure; it was therefore fastened to the bottom end-fitting of the aerobot and would likely be further integrated into this end-fitting in a final mission configuration. The mechanical design centered around accommodating the volumes and masses of the C3 (Command, Control, and Communications) unit, pumps, valves, avionics, and batteries. A 3D-printed chassis was used to minimize structural mass and an aluminum honeycomb sandwich panel was used to interface with the bottom end-fitting of the balloons. Cables from the BCM to the flight termination poppet and skin temperature sensors ran along the exterior of the balloon for easy engineering access, adhered to the envelope with FEP tape. Figure 5a shows annotated pictures of the BCM; the total mass of the BCM was 19.5 kg.
@@ 2.2.2 Assembly-level Payload Features
Measurements made on the gondola and ground station were saved on-board and retrieved after the flight test. Data collected on-board the BCM were stored on board at the native sampling rate (variable, but $>10$ Hz for most instruments), and relayed back to the ground station through a radio link at 1 Hz for redundancy and flight monitoring. The measurements, commanding, and communication on the three platforms was performed by an identical flight computer stack. This ensured interchangeability and redundancy in the event of any failure. The flight computer stack and avionics diagram are shown in the supplementary material, along with post-flight calibration data of some of the instruments.
@@ 2.3.1 Launch Site Selection
As discussed in Section 1, a low altitude flight in Earth atmosphere serves as a good analog for flights in Venus atmosphere in the target altitude range. This being the first outdoor flights of the prototype Venus aerobot, one of the main criteria in choosing a field site was the need for being able to terminate the flight at any point without risking either damage to the one-of-a-kind prototype by ground scrub brush or an unsafe overflight of populated areas. The Black Rock desert in Nevada was chosen as it offers many kilometers of unobstructed, flat lakebed for a chase vehicle to follow the balloon with a high likelihood of safe and speedy recovery upon landing on predictable smooth terrain. Bureau of Land Management and Federal Aviation Administration permissions were obtained prior to flight.
@@ 2.3.2 Mission Planning
The other aspect of wind forecast relevant to mission planning is surface winds. Inflating, trimming the buoyancy of the balloon, and recovering the aerobot during high winds is challenging and potentially damaging for the balloon envelopes. On most days of the campaign, the winds were lowest around sunrise. Our goal was to launch with surface winds below 5 mph and recover with surface winds below 8 mph based on expertise at Aerostar. However, a sunrise launch also has significant undesirable thermal implications on the balloon - creating the largest helium and atmospheric temperature variations between launch and ‘cruise’ phases of the flight under solar heating. This translates to a risk of flying to higher-than-desired altitudes or ground recontacts. Keeping these factors in mind, we selected a launch time close to sunrise for both the launches but kept the target free-lift to under 500 g, as ground recontact could be mitigated by the recovery teams.
@@ 2.3.3 Flight Tests
Figure 6 shows four images from the flight tests. Instrument checkouts were carried out a day before each launch. On the day of each launch, given the need to launch at sunrise, setup and operations began close to midnight. After a final integration and checkout of all the sub-systems, the super-pressure balloon was inflated, followed by the zero-pressure balloon, until the desired free-lift was obtained. The winds observed, especially during the first launch, were much higher than expected (in excess of 10 mph) possibly due to topography near the lakebed boundaries. This made the launch operations more challenging and reduced our ability to accurately estimate the free lift of the aerobot at launch. Consequently, in both the launches, the aerobot needed to be brought back to the ground and team members in the chase vehicle had to re-ballast the payload gondola to achieve the desired free-lift condition. In the following sections of the paper, we primarily focus on the phases of the missions after the re-ballasting was done.
Figure 6: Clockwise from top left: equipment and aerobot being laid out on tarp prior to inflation (4:53AM); inflated aerobot prior to release (7:24AM); aerobot in free flight (7:33AM); aerobot being deflated after landing (9:28AM). Photographs from Flight 1.
@@
Figure 7 shows the ground tracks of the aerobot for the two flights. Note that in the second flight, the turn in the heading of the balloon was carried out through altitude adjustment by the ground operations team, to move the aerobot towards the center of the lakebed based on predictions of how the wind varied with altitude.

### v2

@@ Abstract
This paper details a significant milestone towards maturing a buoyant aerial robotic platform, or aerobot, for flight in the Venus clouds. We describe two flights of our subscale altitude-controlled aerobot, fabricated from the materials necessary to survive Venus conditions. During these flights over the Nevada Black Rock desert, the prototype flew at the identical atmospheric densities as 54 to 55 km cloud layer altitudes on Venus. We further describe a first-principle aerobot dynamics model which we validate against the Nevada flight data and subsequently employ to predict the performance of future aerobots on Venus. The aerobot discussed in this paper is under JPL and Aerostar development for an in-situ mission flying multiple circumnavigations of Venus, sampling the chemical and physical properties of the planet’s atmosphere and also remotely sensing surface properties.
©2025 California Institute of Technology. Accepted document version. Published in AIAA Journal of Aircraft (JA) 2025. DOI: https://doi.org/10.2514/1.C038314
@@
$\theta$  = angle from vertical, radians
@@ 1 Introduction
Figure 1: Subscale prototype of JPL’s variable-altitude controlled aerobot, made from Venus-compatible materials, in flight over the Black Rock desert in Nevada, USA. Photograph taken July 2022, courtesy Aerostar. ©2022 California Institute of Technology.
@@ 1.1 Venus Aerobot Background
Balloon experiments have operated on Venus before but none of the proposed capability. The Soviet VeGa lander missions of 1985 included balloon probes with two small 7 kg gondolas that flew for approximately 46 hours in the cloud layer at a nominally constant altitude [1]. The challenge today is in expanding this early demonstrator to a larger and fundamentally more capable buoyant aerial robotic platform (aerobot), thereby enabling the greater science return promised by a dedicated mission. NASA’s Jet Propulsion Lab (JPL) has been developing large aerobots for specifically this purpose for over twenty years, including multiple Discovery proposals [2], carrying a 40-100 kg gondola below the aerobot with science instrumentation consistent with a much more complete characterization of the Venus cloud layer. These aerobots, like the VeGa balloons before, were designed to maintain altitude through the use of "superpressure" - i.e. a high strength envelope inflated to pressure slightly above the ambient atmospheric pressure. Given its constant volume, a superpressure balloon loses buoyancy when displaced upwards and gains buoyancy when displaced downwards, making it altitude stable.
@@
More recently, a NASA study described in Cutts et al. [3] refocused JPL efforts on an aerobot that can instead actively control its altitude over a variable range (Figure 1) instead of staying at a constant altitude. A variable-altitude approach allows the reuse of a single payload over a wide range of conditions in the clouds, thereby more fully characterizing the cloud layer (and providing an even greater science return) for a modest complexity cost. The success of the Google Loon program [4] provided credence to the rapid development of variable altitude aerobots for Earth, and the Cutts et al. [3] NASA study prioritized the adaption of this technology for Venus. A number of variable-altitude Venus design points and their development are described in [5], with targeted altitude range on Venus from 52 to 62 km and a flight duration goal of 117 Earth-days (one Venus day). Such platforms are highly scaleable, with designs able to carry from 15 kg to 230 kg depending on mission ambition.
@@
This paper describes a significant milestone in the technology advancement of this Venus aerial plaform: JPL’s first outdoor flights of a prototype aerobot, made from Venus-compatible materials, at atmospheric densities within the proposed flight envelope for Venus. Table 2 describes the US Standard Atmosphere [6] that approximates the test flight conditions in comparison to the Venus atmosphere [7]. Balloon flight dynamics are principally driven by atmospheric density, meaning that this flight recreates a relevant environment for flight dynamics in Venus’s cloud layer from 54-55 km above the surface. The targeted Venus temperatures are roughly $15^{\circ}$C greater than those in Nevada however, so models must be employed to take this and other environmental differences into account. Accordingly, a key objective for this flight was to provide validation data for our flight dynamics model FLOATS (FLight Operations and Aerobot Trajectory Simulator), which we subsequently use to predict mission trajectories and operations on Venus. While flights at higher altitudes (up to 62 km equivalent density) are possible with our platform, balloons must trade altitude for payload mass for a given envelope volume, and a complete characterization of the balloon dynamics for FLOATS required more payload mass than the full altitude range could support on a prototype of this scale.
Table 2: Atmospheric equivalencies between Venus balloon altitudes and test flight altitudes (bolded rows) in Nevada, USA. Note exact density match, although Venus is slightly warmer.
@@
(incl. $20^{\circ}$C local offset condition) [6]
@@ 1.2 Programmatic Context
The variable-altitude aerobot was accordingly adopted into two OWL-support mission concept studies for Venus [10, 11], and Cutts et al. [12] describe a variety of smaller missions that could be performed with a variable-altitude aerobot, including concepts for New Frontiers and Discovery proposals. The prototype flown in the flight tests described in this paper is approximately one-third diameter scale of the favored design point of these previous mission studies.
@@ 1.3 Paper Outline
A video of the test flights in this paper can be found at the media release accompanying the experiment at https://www.jpl.nasa.gov/news/jpls-venus-aerial-robotic-balloon-prototype-aces-test-flights.
@@ 2.1 Aerobot Prototype Specifications
Figure 2: Photo of aerobot at takeoff, with the two balloon chambers (SP and ZP) labeled. Both balloons are made from the materials needed for a Venus mission, and helium is pumped/vented between balloons for altitude control. Photo courtesy NASA/JPL-Caltech, ©2022 California Institute of Technology.
@@
Figure 3 illustrates the design of the two balloon envelopes. The outer ZP is a bilaminate of Teflon® fluorinated ethylene propylene (FEP) and Kapton® polyimide, with two metalization layers and a bonding adhesive. The Teflon® offers excellent resistance to the clouds of sulfuric acid aerosols during Venus environment evaluations [18], while the Kapton® provides a helium gas barrier and the tensile strength for carrying the payload. The silver metalization acts as a second surface mirror (with a measured solar absorptivity $\alpha=0.08$ and infrared emissivity $\epsilon=0.52$) to reduce solar heating of the helium, and second metalization layer adds extra protection to reduce the effect of helium diffusion through the material. While this specific material has not gone to permeability testing, a prior material of similar metalization construction showed < 5 cm3/m2-day as measured at room temperature [19] consistent with the desired lifetime on Venus. The bilaminate design further adds resistance to pinholing from manufacturing defects, as the chances of two defects aligning between the layers is extremely small. The ZP balloon is constructed as a 60-degree sphere-cone-sphere with upper diameter of 5 meters and lower diameter of 2.5 meters.
Figure 3: A: Material stackup of the zero-pressure (ZP) balloon bilaminate. B: Two layers of the super-pressure (SP) balloon. C: Photograph of SP balloon. Modified with permission from Izraelevitz et al. [13], ©2021 California Institute of Technology.
The inner SP balloon is fabricated as a 2.5 m diameter sphere, also two-layered, with a Vectran® fabric (200 denier 2x2 basket weave at 76 threads-per-inch) to provide pressurization strength and a heat-sealed thermoplastic urethane bladder for gas retention. The two layers are indexed together with attachment tabs to keep relative alignment. While the ZP material is metalized, the urethane need not be. Helium diffusion through the SP balloon is not lost to the atmosphere but goes into the ZP balloon and hence the only effect is to slightly increase the pumping load, which was found to be easily accommodated by the pumping rate during assembly-level testing [19]. The SP balloon was proof-tested to 14 kPa before the flight, and an identical SP balloon was taken to 61 kPa to provide confidence in the planned pressurization margin utilized in the flight tests.
@@ 2.2 Instrumentation, Gas Transfer, and Communications Payload
For accurate measurement of the flight dynamics of the aerobot, we designed and implemented an instrumentation system (Figure 4) that could not only gather information about the 6-DOF dynamics of the aerobot, but also on its thermal state and driving inputs from the environment such as atmospheric temperature, pressure, wind, and solar radiation. In addition, our instrumentation package coordinated the gas transfer between the two aerobot chambers, executed flight termination processes, and sent back real-time telemetry. This payload module was fabricated using hobby-grade avionics and commercial-of-the-shelf instrumentation for ease of development timeline, and would be replaced by spaceflight-grade systems for a Venus mission. This payload is a second iteration of what was flown in our previous indoor flights in Tillamook Airship Hangar [13] with additions for wind measurements, solar forcing, long distance commanding, and Federal Aviation Administration (FAA) requirements. Components are summarized in Table 3. Some sensors (such as the temperature sensors) are sampled much faster than the expected time scale of the dynamics of the measured quantities. This was done primarily for two reasons – first, the presence of a large number of data points to average reduces the noise in the measurement, and second, a fast sample rate was well within system capabilities and allowed for an abundance of caution against any aliasing in a untested dynamical environment.
@@
[…]
Texas Instruments LM35-CA/Z Skin temperature sensor x4 and atmospheric temperature x1 BCM 20
@@
[…]
Windmaster Pro Ultrasonic anemometer for relative wind and atmos. temperature Gondola 20
@@
Inspeed Cup Anemometer Groundspeed winds (display only, not logged) Boom above transport trailer 1
@@ 2.2.1 Payload Platforms
    Buoyancy Control Module (BCM): The BCM contained the entire gas transfer system, including a 12.7 mm (1/2 inch) inner diameter bypass tube through the entire vertical extent of the SP balloon to transfer gas into the ZP balloon. The BCM system needed to be as close to the balloon chambers as possible to reduce additional flow path length and associated loss of pressure; it was therefore fastened to the bottom end-fitting of the aerobot and would likely be further integrated into this end-fitting in a final mission configuration. The mechanical design centered around accommodating the volumes and masses of the C3 (Command, Control, and Communications) unit, pumps, valves, avionics, and batteries. A 3D-printed chassis was used to minimize structural mass and an aluminum honeycomb sandwich panel was used to interface with the bottom end-fitting of the balloons. Cables from the BCM to the flight termination poppet and skin temperature sensors ran along the exterior of the balloon for easy access, adhered to the envelope with FEP tape and covered with ZP material scraps to insulate against solar radiation. Figure 5a shows annotated pictures of the BCM; the total mass of the BCM was 19.5 kg.
@@ 2.2.2 Assembly-level Payload Features
Measurements made on the gondola and ground station were saved on-board and retrieved after the flight test. Data collected on-board the BCM were stored on board at the native sampling rate (variable, but >10>10 Hz for most instruments), and relayed back to the ground station through a radio link at 1 Hz for redundancy and flight monitoring. The measurements, commanding, and communication on the three platforms was performed by an identical flight computer stack. This ensured interchangeability and redundancy in the event of any failure. The flight computer stack and avionics diagram are shown in the supplementary material, along with post-flight calibration data of some of the instruments.
@@ 2.3 Flight Experiment Summary
The field team included nine attendees: five JPL (Jacob Izraelevitz, Siddharth Krishnamoorthy, Ashish Goel, Gerald Walsh, Michael Pauken) and four Aerostar (Tim Lachenmeier, Caleb Turner, Allen Dial, Richard Bauer).
@@ 2.3.1 Launch Site Selection
As discussed in Section 1, a low altitude flight in Earth atmosphere serves as a good analog for flights in Venus atmosphere in the target altitude range. This being the first outdoor flights of the prototype Venus aerobot, one of the main criteria in choosing a field site was the need for being able to terminate the flight at any point without risking either damage to the one-of-a-kind prototype by ground scrub brush or an unsafe overflight of populated areas. The Black Rock desert in Nevada was chosen as it offers many kilometers of unobstructed, flat lakebed for a chase vehicle to follow the balloon with a high likelihood of safe and speedy recovery upon landing on predictable smooth terrain. Bureau of Land Management permissions were obtained prior to flight for use of the lakebed, and flight vehicle parameters were filed in advance as a Certificate of Waiver or Authorization (COA) with the FAA. FAA was again notified by phone before and after the flights, and tracked live by the Oakland Air Route Traffic Control Center (ZOA).
@@ 2.3.2 Mission Planning
The other aspect of wind forecast relevant to mission planning is surface winds. Inflating, trimming the buoyancy of the balloon, and recovering the aerobot during high winds is challenging and potentially damaging for the balloon envelopes. On most days of the campaign, the winds were lowest around sunrise. Our goal was to launch with surface winds below 2 m/s and recover with surface winds below 3.5 m/s based on expertise at Aerostar. However, a morning launch also has significant undesirable thermal implications on the balloon - creating the largest helium and atmospheric temperature variations between launch and ‘cruise’ phases of the flight under solar heating. This translates to a risk of flying to higher-than-desired altitudes or ground recontacts. Keeping these factors in mind, we selected a launch time close to sunrise for both the launches but kept the target free-lift (buoyancy above neutral) to under 500 g, as ground recontact could be mitigated by the recovery teams.
Fourteen T-bottles of helium (8.5m3 per bottle) were brought to Nevada for the two flights, with the expectation of six bottles per flight (and one tank of margin) to meet target the desired lift condition. All equipment was brought to and from the lakebed each morning from accommodations in Gerlach in a array of sport utility vehicles, a towed groundstation trailer, and a recreation vehicle for crew breaks.
@@ 2.3.3 Flight Tests
Figure 6 shows four images from the flight tests. Instrument checkouts were carried out a day before each launch. On the day of each launch, given the need to launch at sunrise, setup and operations began close to midnight. After a final integration and checkout of all the sub-systems, the SP balloon was inflated, followed by the ZP balloon, until the desired free-lift was obtained. This free-lift was measured and trimmed directly from the balloon buoyancy in the minutes before launch, as estimating indirectly from helium input (i.e. tank pressure or a mass flow meter) accumulates significant error with the large gas volume compared to the desired lift and is sensitive to the change in environmental temperatures over the hours between fill and launch.
The winds observed from the cup anemometer, especially during the first launch, were much higher than expected (in excess of 4 m/s) possibly due to topography near the lakebed boundaries. This made the launch operations more challenging and reduced our ability to accurately estimate the free-lift of the aerobot. Consequently, in both the launches, the aerobot needed to be brought back to the ground, and team members in the chase vehicle had to re-ballast the payload gondola to achieve the desired free-lift condition. In the following sections of the paper, we primarily focus on the phases of the missions after the re-ballasting was done.
Figure 6: Clockwise from top left: equipment and aerobot being laid out on tarp prior to inflation (4:53AM); inflated aerobot prior to release (7:24AM); aerobot in free flight (7:33AM); aerobot being deflated after landing (9:28AM). Photographs from Flight 1, courtesy NASA/JPL-Caltech. ©2022 California Institute of Technology.
@@
Figure 7 shows the ground tracks of the aerobot for the two flights. Note that in the second flight, the turn in the heading of the balloon was carried out through altitude adjustment by the ground operations team, to move the aerobot back towards the center of the lakebed based on predictions of how the wind varied with altitude. The wind predictions to inform this turn were obtained in real time by the pilot via GFS/ECMWF models from Windy.com [22].